What are the x-intercepts of the graph of y=(2x+3)(x-5)?
A) -3/2 and 5
B) -3/2 and -5
C) 3/2 and 5
D) 3/2 and -5

Answer by stanbon(75887)

What are the x-intercepts of the graph of y=(2x+3)(x-5)?
----
All x-intercepts have y = 0.
---
Your Problem:
2x+3 = 0 or x-5 = 0
x = -3/2 or x = 5
-------------------------
Cheers,
Stan H.
=================
A) -3/2 and 5
B) -3/2 and -5
C) 3/2 and 5
D) 3/2 and -5